OpenStudy (anonymous):

Part 1: Find the polynomial f(x) that has the roots of –3, 5 of multiplicity 2. (4 points) Part 2: Explain how you would verify the zeros of f(x). (4 points)

OpenStudy (whpalmer4):

If you have a collection of roots \(r_1, r_2,...,r_n\) and want to construct the polynomial, write a series of product terms \[P(x) = (x-r_1)(x-r_2)...(x-r_n)\]If a root has multiplicity > 1, you need to repeat it that many times in the product. You'll probably be expected to expand that product. To verify the zeros, one way would be to check that the polynomial will have a value of 0 at each zero. Another would be to divide the polynomial by \((x-r_n)\) (to verify that \(r_n\) is a zero) — the division will have a remainder of 0 is \(r_n\) is a zero of the polynomial.

OpenStudy (whpalmer4):

Another way, of course, would be to graph the polynomial...

OpenStudy (whpalmer4):

The value of \(x\) where the curve touches or crosses the x-axis will be zeros of the polynomial.

OpenStudy (anonymous):

so its (x+3) (x-5) (x-2) ? @whpalmer4

OpenStudy (whpalmer4):

Nope, reread what I said about multiplicity.

OpenStudy (anonymous):

dont you foil now ? @whpalmer4

OpenStudy (whpalmer4):

Also, the problem is poorly worded: there are infinitely many polynomials with those zeros, because you can have a constant multiplier in front of all of the product terms without changing the zeros. It is inappropriate to speak of "the polynomial with zeros" unless there is a specified point which the graph of the polynomial goes through, which then uniquely identifies that constant multiplier. The problem should have said "find A polynomial ..."

OpenStudy (whpalmer4):

Could you make me a list of all of the roots?

OpenStudy (anonymous):

i dont understand

OpenStudy (whpalmer4):

you need a list of all of the zeros (aka roots) to do this. what are they?

OpenStudy (whpalmer4):

" polynomial f(x) that has the roots of –3, 5 of multiplicity 2" what does that mean to you?

OpenStudy (anonymous):

it means they made it equal to zero. so it used to be (x+3) (x-5) (x-2)

OpenStudy (whpalmer4):

nope. make me a list of roots. x = ?

OpenStudy (whpalmer4):

"roots of -3, 5 of multiplicity 2" - what does that mean?

OpenStudy (anonymous):

x=-3 x=5 x=2

OpenStudy (whpalmer4):

okay, what does multiplicity mean?

OpenStudy (anonymous):

i dont know

OpenStudy (whpalmer4):

hint: "multiplicity 2" does not necessarily mean that x=2 is a root/zero!

OpenStudy (anonymous):

multiply ?

OpenStudy (whpalmer4):

okay, here's a tip: if you don't understand all the words in the problem, you probably aren't going to get the right answer!

OpenStudy (whpalmer4):

multiplicity means the number of times a root/zero is repeated. for example, \[y=x^2+2x + 1\] has a root of \(x=1\) with multiplicity 2 because it can be factored as \(x^2+2x+1 = (x+1)(x+1)\)

OpenStudy (whpalmer4):

If a root has an odd multiplicity, the curve crosses the x-axis at the zero/root. If the root has an even multiplicity, the curve either dips down to or climbs up to and touches the x-axis at the root, then retreats. For example, \(y = x^2\) has multiplicity 2, and as it is a parabola, the curve drops to the x-axis at x = 0, makes contact, and then climbs away. \(y = x\) on the other hand has multiplicity 1 for the root at \(x = 0\) so the curve/line crosses through the x-axis at x = 0.

OpenStudy (anonymous):

alrighty soooooo lol

OpenStudy (whpalmer4):

So, here's a spot where the question a bit ambiguous in my opinion: "roots of –3, 5 of multiplicity 2" does that mean roots of -3, 5, 5 or -3, -3, 5, 5 Probably it means the first one, but were I doing the problem, I would state outright that that was my interpretation of the problem statement, in case it turned out to be the other that was expected.

OpenStudy (whpalmer4):

So, you need to make a choice: -3, 5, 5 or -3, -3, 5, 5 Then, write out the polynomial: \[(x+3)(x-5)(x-5) = (x+3)(x-5)^2 = ...\] or \[(x+3)(x+3)(x-5)(x-5) = (x+3)^2(x-5)^2 = ...\]

OpenStudy (whpalmer4):

when you expand the polynomial, the highest exponent on x will be the number of product terms you have.